A string is wrapped over the edge of a uniform disc and the free end is fixed with the ceiling. The disc moves down, unwinding the string. Find the downward acceleration of the disc.

Grade:9

1 Answers

Akash Kumar Dutta
11 years ago

Dear Vrushabh,
T.R=I.w
M(g-a).R=M(R^2).a/2R
(g-a)=a/2
g=3/2a
a=2/3g (ANS)
Regards.